The number of emails you sent can be modelled by the equation
u = 20×(1.8)^e,
where u is the total number of emails sent, and e is the number of
months since you started sending emails, where 0≤ e ≤ 8.
1 - find the month in which the exponential model predicts that the total
number of emails sent will first reach 500?
2 - Write down the value of the scale factor for the exponential
model. Use this to find the percentage increase in the total
number of emails sent?

Answer by Alan3354(69443)

The number of emails you sent can be modelled by the equation
u = 20×(1.8)^e,
where u is the total number of emails sent, and e is the number of
months since you started sending emails, where 0≤ e ≤ 8.
1 - find the month in which the exponential model predicts that the total
number of emails sent will first reach 500?
2 - Write down the value of the scale factor for the exponential
model. Use this to find the percentage increase in the total
number of emails sent?
------
I'll use t for the # of months.
Also, use * (Shift 8) for multiply. x is often used elsewhere.
-------
u = 20*(1.8)^t
---
1 - find the month in which the exponential model predicts that the total
number of emails sent will first reach 500?
500 = 20*(1.8)^t
25 = 1.8^t
log(25) = t*log(1.8)
t = log(25)/log(1.8)
t =~ 5.48 months
--> the 5th month
==============================
2 - Write down the value of the scale factor for the exponential
model. Use this to find the percentage increase in the total
number of emails sent?
Increase from when to when?